OpenStudy (anonymous):

Which equation represents the line that passes through (–6, 7) and (–3, 6)?

OpenStudy (anonymous):

y = –x + 9 y = –x + 5 y = –3x – 11y y = –3x + 25

OpenStudy (anonymous):

first we find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (-6,7)....x1 = -6 and y1 = 7 (-3,6)...x2 = -3 and y2 = 6 now we sub slope(m) = (6 - 7) / (-3 - (-6) slope(m) = -1/ (-3 + 6) slope(m) = -1/3 now we use y = mx + b slope(m) = -1/3 use either of your points (-6,7)...x = -6 and y = 7 now we sub 7 = -1/3(-6) + b 7 = 2 + b 7 - 2 = b 5 = b so our equation is : y = -1/3x + 5 I am getting something totally different then your answer choices

OpenStudy (anonymous):

are you sure your problem is not asking for a perpendicular line that passes through those points ?